OpenStudy (majikdusty):

Find the Values for X so that the series converges. The series is from n=1 to Inf of 10/x^n What I did was remove the 2 from the fraction so I got 10(1/x)^n. I took the abs(1/x)<1 because geometric series must be below 1 to converge. And the answer i received was (-1,0)U(0,1) However this answer is wrong. I am sure I did something wrong, but i just do not know what.

OpenStudy (amistre64):

take a limit or the ratio (n+1)/n

OpenStudy (amistre64):

\[\lim\frac{10x^{-n-1}}{10x^{-n}}\] \[\lim\frac{x^{-n-1}}{x^{-n}}\] \[\lim x^{-n+n+1}=|x|\] \[|x|<1~:~-1<x<1\]

OpenStudy (majikdusty):

Hmm I'm not quite sure what that ratio thing is. We are currently in Geometric series in the book right now.